# Robust Adaptive Discrete-Time Control Barrier Certificate

### v1

@@ Abstract
This work develops a robust adaptive control strategy for discrete-time systems using Control Barrier Functions (CBFs) to ensure safety under parametric model uncertainty and disturbances. A key contribution of this work is establishing a barrier function certificate in discrete time for general online parameter estimation algorithms. This barrier function certificate guarantees positive invariance of the safe set despite disturbances and parametric uncertainty without access to the true system parameters. In addition, real-time implementation and inherent robustness guarantees are provided. Our approach demonstrates that, using the proposed robust adaptive CBF framework, the parameter estimation module can be designed separately from the CBF-based safety filter, simplifying the development of safe adaptive controllers for discrete-time systems. The resulting safety filter guarantees that the system remains within the safe set while adapting to model uncertainties, making it a promising strategy for real-world applications involving discrete-time safety-critical systems.
@@ 1 Introduction
Ensuring the safety of dynamical systems is a fundamental requirement in modern control applications, particularly in domains where violations of safety constraints can lead to undesirable outcomes [4], [36]. Safety-critical control focuses on the synthesis of controllers guaranteeing that the system remains within a safe set. Among various approaches (e.g., Hamilton-Jacobi reachability analysis [5], penalty functions [33], and Model Predictive Control (MPC) [20], [30]), Control Barrier Functions (CBFs) have emerged as a powerful and systematic tool to enforce safety for both continuous-time [4] and discrete-time systems [2]. CBF-based control methods allow the formal encoding of safety constraints, and they have found wide-ranging applications, including autonomous driving [8], robotics [16], and power systems [41], demonstrating their versatility and practical relevance in safety-critical domains.
@@
Despite the recent surge in learning-based approaches for safety-critical control [35], model-based methods using CBFs remain essential due to their analytical tractability, real-time implementability, and formal safety guarantees [4], [43]. These approaches allow for rigorous controller design based on dynamical models [2], [4], [46], making them highly attractive for safety-critical applications. However, a fundamental challenge in deploying model-based CBFs is the presence of model uncertainty, which can compromise safety if not properly addressed [43]. To mitigate this, various extensions have been developed that incorporate robust and/or adaptive mechanisms against disturbances [12], [13], [18], [48] or parametric uncertainty [24], [25], [34], [38], [44]. Specifically, for safe control under disturbances, verifiable certificates have been established to ensure safety with high probability for unbounded stochastic disturbances [12], [13] and guaranteed safety for bounded disturbances [18], [48]. On the other hand, recent results on adaptive safety under parametric model mismatch have been obtained using CBFs with online parameter updates, employing either Lyapunov-based methods [24], [25], [38], [44] or batched least-squares methods [34]. For an in-depth overview, we refer the reader to [43].
Although recent developments are promising, to the best of the authors’ knowledge, existing results on adaptive safety for handling parametric uncertainty are limited to continuous-time settings [24], [25], [34], [38], [44], while the discrete-time adaptive case remains unexplored. Given the widespread use of digital controllers and discrete-time formulations in practice—such as in MPC and reinforcement learning [42]—it is critical to develop adaptive safety control within a Discrete-Time CBF (DT-CBF) framework to bridge this gap. Moreover, such discrete-time frameworks are more compatible with real-world applications with discrete-time sensing and actuation [12], [32].
Since time derivatives are not available in discrete time, the main challenge lies in establishing all results using difference equations and inequalities. Moreover, a causal synthesis of the online safe controller must be ensured such that no future data is required for computing the control input., which is an important aspect that is missing when continuous-time models are used [25], [34]. On the other hand, Lyapunov-based adaptation laws often lack guarantees on the boundedness of the parameter estimates [44] or convergence to the true parameters [25], making estimation-based approaches more appealing [10], [44]. Since many simple and effective estimation or identification methods (e.g., least-squares techniques) require access to sampled data [23], a DT-CBF framework is naturally better suited for leveraging these methods. In this paper, we address this theoretical gap by developing a DT-CBF framework for adaptive safety.
@@
Our approach extends model-based safe adaptive control to a broader class of systems—particularly those implemented in embedded or sampled-data settings—and lays the groundwork for online safe control in real-world systems with uncertain models. Specifically, the contributions of this work can be summarized as follows:
  • We develop a robust adaptive Control Barrier Certificate (CBC) for input-affine discrete-time systems subject to both bounded disturbances and parametric model mismatch. To the best of our knowledge, this is the first result on adaptive safety using DT-CBF.
@@
  • The proposed certificate accommodates general online adaptation laws that are decoupled from the CBF-based controller synthesis, enabling a modular control architecture in the spirit of [9]. This new insight—similar to the separation principle in the control and estimation for linear systems [39]—has not been revealed in existing results [25], [34].
@@
  • The inherent robustness111 In this paper, inherent robustness refers to steering the state towards a safe set when the current state is unsafe [46, Section 2.2], which should not be confused with robustness against model uncertainty [15], [25]. of the proposed robust adaptive control approach using DT-CBF is rigorously analyzed, extending existing robustness guarantees for continuous-time systems with time-invariant control barrier certificates [46].
@@
The rest of the paper is organized as follows. Section 2 provides preliminaries. Section 3 introduces the robust adaptive CBC using DT-CBFs. A simple online estimation algorithm based on Recursive Least Squares (RLS) is briefly given in Section 4. Inherent robustness guarantees are discussed in Section 5. A simulation example is presented in Section 6, and Section 7 concludes the paper and outlines future research directions.
@@ 2 Preliminaries
This section introduces robust DT-CBFs, which constitute a natural and modest extension of continuous-time CBFs to discrete-time systems with bounded disturbances, building on prior work on continuous-time systems with bounded disturbances [18] and discrete-time systems without disturbances [2].
@@ 2.1 Notation
The sets of real and natural numbers are denoted, respectively, by $\mathbb{R}$ and $\mathbb{N}$. The set of extended natural numbers is $\overline{\mathbb{N}}=\mathbb{N}\cup\{+\infty\}$, and $\mathbb{I}_{[a:b]}:=\overline{\mathbb{N}}\cap[a,b]$. The matrix (vector) $p$-norm is denoted by $\|\cdot\|_{p}$ with $p\in\{1,2\}$, and $\|\cdot\|$ represents the $2$-norm by default. The $i$-th element of a vector $x$ is denoted by $x[i]$. The zero and one vectors of size $n$ are denoted by $\mathbf{0}_{n}$ and $\mathbf{1}_{n}$, respectively. A continuous function $\alpha:\mathbb{R}\to\mathbb{R}$ is said to be an extended class $\mathcal{K}_{\infty}$ function (denoted as $\alpha\in\mathcal{K}_{\mathrm{e},\infty}$) if $\alpha(0)=0$, $\alpha$ is strictly monotonically increasing, $\lim_{r\to-\infty}\alpha(r)=$-$\infty$, and $\lim_{r\to+\infty}\alpha(r)=+\infty$ [38, Definition 7]. The class of functions $\beta:\mathbb{R}\to\mathbb{R}$ that are sublinear w.r.t. identity (i.e., $\beta(r)<r$ for all $r>0$) is denoted by $\mathcal{C}_{\mathrm{SL}}$. The projection operator $\Pi_{\mathcal{X}}(\cdot)$ and distance operator $\mathrm{dist}_{\mathcal{X}}(\cdot)$ w.r.t. a set $\mathcal{X}\subseteq\mathbb{R}^{n}$ are defined, respectively, as $\Pi_{\mathcal{X}}(\bar{x})=\arg\min_{x\in\mathcal{X}}\|x-\bar{x}\|$ and $\mathrm{dist}_{\mathcal{X}}(\bar{x})=\min_{x\in\mathcal{X}}\|x-\bar{x}\|$ for any $\bar{x}\in\mathbb{R}^{n}$. The minimum eigenvalue of a symmetric positive definite matrix $\Gamma$ is denoted by $\lambda_{\min}(\Gamma)$.
@@ 2.2 System description
$=f_{\mathrm{d}}(x_{t})-[\bm{\phi}(x_{t})]^{\top}\theta+g(x_{t})u_{t}+w_{t},\vskip-5.69046pt$ (1)
where $x_{t}\in\mathcal{X}\subset\mathbb{R}^{n}$ and $u_{t}\in\mathcal{U}\subseteq\mathbb{R}^{m}$ are the state and input at time step $t$ with $\mathcal{X}$ being compact and $\mathcal{U}$ being closed, $f_{\mathrm{d}}:\mathcal{X}\to\mathbb{R}^{n}$ describes the nominal drift [^2] Although we consider discrete-time systems, we adopt the term “drift” from the “drift vector field” used in nonlinear control of input-affine continuous-time systems [29]., $g:\mathbb{R}^{n}\to\mathbb{R}^{n\times m}$ characterizes the input coupling, and $\bm{\phi}(x_{t})]^{\top}\theta$ represents the additional drift term with $\bm{\phi}:\mathbb{R}^{n}\to\mathbb{R}^{q\times n}$ being a continuous known kernel (i.e., regressor) and $\theta\in\Theta$ being an unknown parameter. The considered model class in (2.2) can be used in many real-world applications, e.g., inverted pendulum [27], throttle control [37] and Adaptive Cruise Control (ACC) [38] in automotive systems. In addition, the disturbance $w_{t}$ lies in a known polytope given by
@@
where $H_{w}\in\mathbb{R}^{n_{w}\times n}$ and $n_{w}$ is the number of linear inequalities that define $\mathcal{W}$. It is clear that $\mathbf{0}_{n}$ lies in the interior of $\mathcal{W}$, and we further define $\overline{w}:=\max_{w\in\mathcal{W}}\|w\|$. The true parameter $\theta^{\ast}\in\Theta$ is a constant and the prior known parameter set $\Theta$ is a polytope given by
$\Theta=\{\theta\in\mathbb{R}^{p}\mid H_{\theta}\theta\leq h_{\theta}\},\vskip-5.69046pt$ (3)
where $H_{\theta}\in\mathbb{R}^{n_{\theta}\times q}$ and $h_{\theta}\in\mathbb{R}^{n_{\theta}}$. Given an estimate $\hat{\theta}\in\Theta$, its associated estimation error is defined by
$e_{\theta}:=\hat{\theta}-\theta^{\ast}.\vskip-5.69046pt$ (4)
The error $e_{\theta}$ cannot be computed exactly since $\theta^{\ast}$ is unknown. However, for a specific estimate $\hat{\theta}$, an upper bound of the estimation error can be computed by solving the following non-convex optimization problem:
@@
and it holds that [^3] Given $x\in\mathbb{R}^{n}$, $\|x\|_{2}\leq\|x\|_{1}$ [17]. $\|e_{\theta}\|\leq\varepsilon_{\theta}(2)\leq\varepsilon_{\theta}(1)$. Following (2.2), the true system under disturbances is given by
$x_{t+1}=f(x_{t},u_{t};\theta^{\ast})+w_{t},\vskip-5.69046pt$ (6)
assuming access to the true parameter $\theta^{\ast}$.
@@
The controller design for the model (2.2) does not require the smoothness of the kernel $\bm{\phi}(\cdot)$. Therefore, with special non-smooth kernels, the model (2.2) can also describe PWA dynamics [28], [42], and our results can also be used for safety-critical control of PWA systems.
2.3 Robust safety certificate
@@
A set $\mathcal{X}_{\mathrm{s}}\subseteq\mathcal{X}$ is positive invariant for (6) if $x_{0}\in\mathcal{X}_{\mathrm{s}}$ implies that $x_{t}\in\mathcal{X}_{\mathrm{s}}$ for all $t\in\mathbb{N}$ and all $w_{t}\in\mathcal{W}$. In this case, the system (6) is robustly safe w.r.t. $\mathcal{X}_{\mathrm{s}}$, and $\mathcal{X}_{\mathrm{s}}$ is called a safe set.
Characterizing safe sets using CBFs is useful for safe controller design [3], [4]. For discrete-time systems, DT-CBFs are powerful tools to achieve this objective [2], [13]. Specifically, given a continuous function $B:\mathcal{X}\to\mathbb{R}$, its induced safe set is defined using the 0-superlevel set as
$\mathcal{S}_{B}:=\{x\in\mathcal{X}\mid B(x)\geq 0\},$ (7)
and its boundary $\partial\mathcal{S}_{B}$ and interior $\mathrm{int}(\mathcal{S}_{B})$ are defined, respectively, by
$\partial\mathcal{S}_{B}$ $:=\{x\in\mathcal{X}\mid B(x)=0\},$ (8a)
$\mathrm{int}(\mathcal{S}_{B})$ $:=\{x\in\mathcal{X}\mid B(x)>0\}.\vskip-5.69046pt$ (8b)
In addition, to certify safety in the discrete-time setting, an additional continuity assumption is required, which is given as follows:
@@
The barrier function $B:\mathcal{X}\to\mathbb{R}$ is locally Lipschitz continuous on the state space $\mathcal{X}$, i.e.,
$|B(x^{\prime})-B(x^{\prime\prime})|\leq L_{B}\|x^{\prime}-x^{\prime\prime}\|,\quad\forall x^{\prime},x^{\prime\prime}\in\mathcal{X},\vskip-5.69046pt$ (9)
where $L_{B}\geq 0$ is the Lipschitz constant.
Assumption 1 has been used in the literature on discrete-time CBFs [27], [48], and given that the commonly used CBFs are continuously differentiable [3], [4], Lipschitz continuity directly follows given that $\mathcal{X}$ is compact. Besides, the CBFs used in many safe control applications are Lipschitz continuous [1], [25], [38], [48]. The function $B$ is a DT-CBF if it satisfies additional conditions w.r.t. the system (6). The following definition is adopted from [2] and [13] with modifications that suit discrete-time models with bounded disturbances.
Definition 2.
A continuous function $B:\mathcal{X}\to\mathbb{R}$ is a robust DT-CBF for (6) on $\mathcal{S}_{B}$ if there exists $\alpha\in\mathcal{K}_{\mathrm{e},\infty}\cap\mathcal{C}_{\mathrm{SL}}$ such that for all $x\in\mathcal{X}$, there exists $u\in\mathcal{U}$ satisfying
$B(f(x,u;\theta^{\ast})+w)-B(x)-L_{B}\overline{w}\geq-\alpha\left(B(x)\right),\vskip-5.69046pt$ (10)
@@
The barrier certificate (10) also defines the set of robust safe control inputs for $\mathcal{S}_{B}$ as
@@
where $\theta^{\ast}$ emphasizes that characterizing $\mathcal{U}_{\mathrm{rs}}(x;\theta^{\ast})$ requires knowing the true parameter. The following theorem is a standard result on CBF-based safety guarantees, and the proof is omitted for brevity since it follows the same procedure as in [2].

### v2

@@ Abstract
This work develops a robust adaptive control strategy for discrete-time systems using Control Barrier Functions (CBFs) to ensure safety under parametric model uncertainty and disturbances. A key contribution of this work is establishing a barrier function certificate in discrete time for general online parameter estimation algorithms. This barrier function certificate guarantees positive invariance of the safe set despite disturbances and parametric uncertainty without access to the true system parameters. In addition, real-time implementation and inherent robustness guarantees are provided. The proposed robust adaptive safe control framework demonstrates that the parameter estimation module can be designed separately from the CBF-based safety filter, simplifying the development of safe adaptive controllers for discrete-time systems. The resulting safe control approach guarantees that the system remains within the safe set while adapting to model uncertainties, making it a promising strategy for discrete-time safety-critical systems.
@@ 1 Introduction
Safety-critical control aims to guarantee that the system state remains within a safe set, i.e., constraint satisfaction, which is crucial in domains where unsafe system behaviors can lead to undesirable outcomes [3]. Among various approaches (e.g., Hamilton-Jacobi reachability analysis [35] and Model Predictive Control (MPC) [21], [32]), Control Barrier Functions (CBFs) becomes a powerful and systematic tool for both continuous-time [3] and discrete-time systems [2]. CBFs allow formal encoding of safety constraints as control invariant sets [4], and they can be integrated on top of traditional control methods (e.g., control Lyapunov functions [3], [19]) and learning-based approaches (e.g., reinforcement learning [9]) that may not inherently offer safety guarantees. CBFs have many applications, including control of bipedal robots [2], Adaptive Cruise Control (ACC) [3], and vehicle engine control [38], demonstrating its versatility and practical relevance in safety-critical domains.
@@
Despite the recent surge in learning-based approaches in CBFs [37], model-based methods remain essential due to their analytical tractability and formal safety guarantees [3], [40]. Typically, these approaches support modular design through the explicit use of dynamical models [2], [3], [43]. A standard approach is to encode safety constraints through a CBF-based inequality condition on the input, and then to incorporate this condition within a safety filter [12] that minimizes the adjustment of a given nominal input that may otherwise lead to unsafe behaviors (i.e., constraint violations). However, a fundamental challenge in model-based methods is model uncertainty, which can compromise safety [40]. Accordingly, various extensions have developed robust and adaptive mechanisms tailored for CBFs to handle disturbances [10], [11], [15], [45] or parametric uncertainty [25], [36], [39], [41]. For disturbances, high-probability safety [11] and guaranteed safety [15] have been established for unbounded and bounded disturbances, respectively. There are also results on adaptive safety focusing on parametric uncertainty, employing either Lyapunov-based methods [25], [39], or batched least-squares methods [36]. For an in-depth overview, we refer the reader to [40].
Although recent progress has been made on robust and adaptive CBFs in continuous time [25], [36], [39], [41], the discrete-time case remains largely unexplored despite its relevance to MPC and reinforcement learning [9]. Since practical systems rely on sampled sensing and actuation, there is a need for adaptive safe control directly based on Discrete-Time CBFs (DT-CBFs) [10], [34]. The main difficulty is that derivative-based arguments are unavailable for DT-CBFs. In addition, a causal controller must compute inputs without using future information, a feature not discussed in continuous-time methods [25], [36]. Besides, Lyapunov adaptation laws lack general guarantees on bounded updates or convergence [41], [25], motivating estimation-based approaches that better align with sampled-data settings [9]. Since standard estimation tools (e.g., least-squares methods) operate on discrete measurements [24], a DT-CBF framework provides a more suitable foundation for adaptive safety.
@@
Our approach extends model-based safe adaptive control to a broader class of systems, particularly those implemented in embedded or sampled-data settings. The contributions of this work are summarized as follows:
  • Robust adaptive safety guarantees using DT-CBFs for input-affine discrete-time systems under bounded disturbances and parametric mismatch are established, and an online input condition is derived to facilitate their deployment in safety filters.
@@
  • The proposed approach accommodates general parameter estimators that are decoupled from the CBF-based controller synthesis, enabling a modular control architecture in the spirit of [8]. This new insight has not been revealed in existing results [25], [36], [39].
@@
  • The inherent robustness [^1] In this paper, inherent robustness refers to steering the state towards safe sets when the current state is unsafe [43, Section 2.2], which should not be confused with robustness against model uncertainty [12], [25].of the proposed safe control approach is rigorously analyzed, extending existing robustness guarantees for continuous-time systems without model uncertainty [43].
@@
It is further noted that the primary contribution of this paper lies in establishing a general robust adaptive safe control framework using DT-CBFs, rather than in designing CBFs, in developing novel parameter estimation algorithms, or in verifying the existence of barrier functions [31]. The rest of the paper is organized as follows. Section 2 provides preliminaries, and Section 3 introduces robust adaptive DT-CBFs. A simple estimation algorithm is given in Section 4. Inherent robustness guarantees are discussed in Section 5. A simulation example is presented in Section 6, and Section 7 concludes the paper and outlines future research directions.
@@ 2 Preliminaries
This section lays the groundwork for the main results. After introducing notations and the system model, we present robust DT-CBFs and clarify how CBFs define safe sets and translate state safety requirements into an input feasibility condition via positive set invariance. In essence, CBFs restrict admissible control inputs to ensure that the state remains within the safe set or is driven back to it if it attempts to leave.
@@ 2.1 Notation
The sets of real and natural numbers are denoted, respectively, by $\mathbb{R}$ and $\mathbb{N}$, $\overline{\mathbb{N}}:=\mathbb{N}\cup\{+\infty\}$, and $\mathbb{I}_{[a:b]}:=\overline{\mathbb{N}}\cap[a,b]$. The matrix (vector) $p$-norm is denoted by $\|\cdot\|_{p}$, and $\|\cdot\|$ represents the $2$-norm by default. The $i$-th element of a vector $x$ is denoted by $x[i]$. The zero vector, one vector, and identity matrix of size $n$ are denoted by $\mathbf{0}_{n}$, $\mathbf{1}_{n}$, and $\mathbf{I}_{n}$, respectively. A continuous function $\alpha:\mathbb{R}\to\mathbb{R}$ is said to be an extended class $\mathcal{K}_{\infty}$ function (denoted as $\alpha\in\mathcal{K}_{\mathrm{e},\infty}$) if $\alpha(0)=0$, $\alpha$ is strictly monotonically increasing, $\lim_{r\to-\infty}\alpha(r)=$-$\infty$, and $\lim_{r\to+\infty}\alpha(r)=+\infty$ [39, Definition 7]. The class of functions $\beta:\mathbb{R}\to\mathbb{R}$ that are sublinear w.r.t. identity (i.e., $\beta(r)<r$ for all $r>0$) is denoted by $\mathcal{C}_{\mathrm{SL}}$. The projection operator $\Pi_{\mathcal{X}}(\cdot)$ and distance operator $\mathrm{dist}_{\mathcal{X}}(\cdot)$ w.r.t. a closed set $\mathcal{X}\subseteq\mathbb{R}^{n}$ are defined, respectively, as $\Pi_{\mathcal{X}}(\bar{x})=\arg\min_{x\in\mathcal{X}}\|x-\bar{x}\|$ and $\mathrm{dist}_{\mathcal{X}}(\bar{x})=\min_{x\in\mathcal{X}}\|x-\bar{x}\|$ for any $\bar{x}\in\mathbb{R}^{n}$. The minimum eigenvalue of a symmetric positive definite matrix $\Gamma$ is denoted by $\lambda_{\min}(\Gamma)$.
@@ 2.2 System description
$=f_{\mathrm{d}}(x_{t})-[\bm{\phi}(x_{t})]^{\top}\theta+g(x_{t})u_{t}+w_{t},\vskip-8.5359pt$ (1)
where $x_{t}\in\mathcal{X}\subset\mathbb{R}^{n}$ and $u_{t}\in\mathcal{U}\subseteq\mathbb{R}^{m}$ are the state and input at time step $t$ with $\mathcal{X}$ being compact and $\mathcal{U}$ being closed, $f_{\mathrm{d}}:\mathcal{X}\to\mathbb{R}^{n}$ describes the nominal drift [^2] Although we consider discrete-time systems, we adopt the term “drift” from the “drift vector field” used in nonlinear control of input-affine continuous-time systems [30].corresponding to known dynamics (e.g., inertia in robotics [11]) and/or $x_{t}$ itself arising from discretization [11], [29], $g:\mathbb{R}^{n}\to\mathbb{R}^{n\times m}$ characterizes the input coupling, and $\bm{\phi}(x_{t})]^{\top}\theta$ represents the uncertain drift term with $\bm{\phi}:\mathbb{R}^{n}\to\mathbb{R}^{q\times n}$ being a continuous known kernel (i.e., regressor) and $\theta\in\Theta$ being an unknown parameter. The uncertain drift can be used to model unknown damping [28] or state-coupled disturbances [39]. A compact state space aligns with bounded safe sets [3], and a closed input space guarantees the existence of feasible solutions to the safety filter optimization problem. These assumptions are practically valid for most engineering systems as they are designed to (or due to physical constraints) lie in a bounded operational regime [2], [3], [25]. The state $x_{t}$ is considered to be fully measured without error, and thus the values $f_{\mathrm{d}}(x_{t})$, $\bm{\phi}(x_{t})$, and $g(x_{t})$ can be obtained accurately. In addition, the disturbance $w_{t}$ lies in a known polytope [^3] A polytope is a bounded polyhedron (i.e., an intersection of a finite set of closed half-spaces) [6, Chapter 2.1].given by
@@
where $H_{w}\in\mathbb{R}^{n_{w}\times n}$ and $n_{w}\geq n$ is the number of linear inequalities that define $\mathcal{W}$. It is clear that $\mathbf{0}_{n}$ lies in the interior of $\mathcal{W}$, and we further define $\overline{w}:=\max_{w\in\mathcal{W}}\|w\|$. The true parameter $\theta^{\ast}\in\Theta$ is a constant and the prior known parameter set $\Theta$ is a polytope given by
$\Theta=\{\theta\in\mathbb{R}^{q}\mid H_{\theta}\theta\leq h_{\theta}\},\vskip-5.69046pt$ (3)
where $H_{\theta}\in\mathbb{R}^{n_{\theta}\times q}$ with $n_{\theta}\geq q$, and $h_{\theta}\in\mathbb{R}^{n_{\theta}}$. Both the disturbance and the model parameter are considered to lie in polytopes, which is motivated by practical considerations, as many disturbances and parameters are naturally bounded and can be modeled by box-constrained sets [1], [3], [25], [45]. Theoretically, polytopes can also inner-approximate arbitrary compact sets, making this a flexible and effective modeling choice. Given an estimate $\hat{\theta}\in\Theta$, its associated estimation error is defined as
$e_{\theta}:=\hat{\theta}-\theta^{\ast}.\vskip-8.5359pt$ (4)
The estimation error $e_{\theta}$ cannot be computed exactly since $\theta^{\ast}$ is unknown. However, for a specific $\hat{\theta}$, an upper bound of its associated estimation error can be computed by solving the following optimization problem:
@@
[…]
$x_{t+1}=f(x_{t},u_{t};\theta^{\ast})+w_{t}.\vskip-5.69046pt$ (6)
The system state evolves according to (6); however, the controller does not have access to the true parameter $\theta^{\ast}$.
@@
The controller design for the model (2.2) does not require the smoothness of the kernel $\bm{\phi}(\cdot)$; with non-smooth kernels, (2.2) can also represent PWA dynamics [29]. The input-affine structure is standard in nonlinear control [30], and linear dependence on the unknown parameters is powerful in kernel modeling [18]. This modeling choice does not trivialize the safe control problem, and it is widely used in practice (e.g., inverted pendulum control [28], throttle control [38], and adaptive cruise control [39]) and enables direct comparison with existing adaptive safe control methods [25], [39].
2.3 Robust safe control
This subsection introduces robust DT-CBFs, which constitute a natural and modest extension of continuous-time CBFs to discrete-time systems with bounded disturbances, building on prior work on continuous-time systems with bounded disturbances [15] and discrete-time systems without disturbances [2].
@@ 2.3 Robust safe control
A set $\mathcal{X}_{\mathrm{s}}\subseteq\mathcal{X}$ is positive invariant for (6) if $x_{0}\in\mathcal{X}_{\mathrm{s}}$ implies that $\forall\{w_{t}\}^{\infty}_{t=0}\in\mathcal{W}^{\infty}$, $x_{t}\in\mathcal{X}_{\mathrm{s}}$ for all $t\in\mathbb{N}$. In this case, the system (6) is robustly safe w.r.t. $\mathcal{X}_{\mathrm{s}}$, and $\mathcal{X}_{\mathrm{s}}$ is called a robust safe set.
Definition 1 considers robustness only against the disturbances, which is valid for the true parameter $\theta^{\ast}$ but may not hold for all parameters $\theta\in\Theta$. Characterizing safe sets using CBFs is useful for safe controller design [3]. For discrete-time systems, DT-CBFs are powerful tools to achieve this objective [2], [11]. Specifically, for a continuous barrier function $B:\mathcal{X}\times\Theta\to\mathbb{R}$, its induced safe set is defined using the 0-superlevel set of $B(\cdot,\theta):\mathcal{X}\to\mathbb{R}$ for all $\theta\in\Theta$ as
$\mathcal{S}_{B,\theta}:=\{x\in\mathcal{X}\mid{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}B(x,\theta)}\geq 0\},\vskip-5.69046pt$ (7)
and its boundary $\partial\mathcal{S}_{B,\theta}$ and interior $\mathrm{int}(\mathcal{S}_{B,\theta})$ are defined, respectively, by $\partial\mathcal{S}_{B,\theta}:=\{x\in\mathcal{X}\mid{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}B(x,\theta)}=0\}$ and $\mathrm{int}(\mathcal{S}_{B,\theta}):=\{x\in\mathcal{X}\mid{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}B(x,\theta)}>0\}$.
Remark 2.
This work adopts barrier functions that depend on model parameter $\theta$ as in [25] and [39] since most barrier functions are model-based [3]. Nonetheless, CBFs that are derived mainly based on the safety requirements without using $\theta$ are also commonly used in many safe control applications, e.g., adaptive cruise control [39], aircraft pitch control [25], and aerial robotics [45].
In addition, a continuity assumption is useful to quantify the perturbation of $B$, which is given as follows:
@@
The function $B:\mathcal{X}\times\Theta\to\mathbb{R}$ is locally Lipschitz continuous on the joint space $\mathcal{X}\times\Theta$, i.e., $\forall x^{\prime},x^{\prime\prime}\in\mathcal{X},\;\theta^{\prime},\theta^{\prime\prime}\in\Theta$,
[…]
with Lipschitz constants $L_{B,x},L_{B,\theta}\geq 0$.
Lipschitz continuity w.r.t. the state $x$ has been used in the literature on discrete-time CBFs [28], [45]. Moreover, when $f(x,u;\cdot):\Theta\to\mathcal{X}$ is continuously differentiable (see the model in (2.2)), the commonly constructed CBFs $B:\mathcal{X}\times\Theta\to\mathbb{R}$ are also continuously differentiable w.r.t. the parameter $\theta$ [25], [39]. Since continuous differentiability on the compact set $\Theta\times\mathcal{X}$ implies Lipschitz continuity, these CBFs are Lipschitz continuous. In addition, many CBFs used in safe-control applications are known to be Lipschitz continuous [2], [25], [39], [45].
Definition 2 (Robust CBF).
Let Assumption 1 hold. The continuous barrier function $B_{\theta^{\ast}}:=B(\cdot,\theta^{\ast})$ is a robust DT-CBF for (6) on $\mathcal{S}_{B,\theta^{\ast}}$ if there exists $\alpha\in\mathcal{K}_{\mathrm{e},\infty}\cap\mathcal{C}_{\mathrm{SL}}$ such that $\forall x\in\mathcal{X}$, $\exists u\in\mathcal{U}$ satisfying
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}B_{\theta^{\ast}}(f(x,u;\theta^{\ast}))\hskip-2.27626pt-\hskip-2.27626ptB_{\theta^{\ast}}(x)}\hskip-2.27626pt-\hskip-2.27626ptL_{B,x}\overline{w}\geq-\alpha\left({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}B_{\theta^{\ast}}(x)}\right),\vskip-5.69046pt$ (8)
@@
The barrier certificate (8) also defines the set of robust safe control inputs for $\mathcal{S}_{B,\theta^{\ast}}$ as
@@
where the subscript $\theta^{\ast}$ emphasizes that characterizing $\mathcal{U}_{\mathrm{rs},\theta^{\ast}}(x)$ requires knowing the true parameter. The following theorem is a standard result on CBF-based safety guarantees, and the proof is omitted for brevity since it follows the same procedure as in [2].